OpenStudy (anonymous):

A pole that is 25 feet tall casts a shadow 65 feet long. At the same time, another pole casts a shadow having a length of 39 feet. What is the height, in feet, of the second pole?

OpenStudy (anonymous):

plllleeease help

OpenStudy (anonymous):

\(\displaystyle \frac{25}{65}=\frac{x}{39}\) Solve for \(x\)

OpenStudy (anonymous):

thanks